OpenStudy (lgbasallote):

LGBARIDDLE: Rearrange the digits and the dots so that their sum is 82 \[\cdot 4 \cdot 5 \cdot 6 \cdot 7 \cdot 8 \cdot 9 \cdot 0\cdot\]

OpenStudy (lgbasallote):

i wonder how long it will take people for this one...

hartnn (hartnn):

what does . represent ? multiplication ?

OpenStudy (anonymous):

What do you mean? The question gives a product involving 0, so the answer is always 0.

OpenStudy (lgbasallote):

\(\cdot\) is a notation for repeating decimals For example \[\cdot 3 = \frac 13\]

hartnn (hartnn):

or can i use it as decimal ?? like 78.6 ?

OpenStudy (anonymous):

Would probably have been better to explain that before, I've never seen that notation before.

OpenStudy (lgbasallote):

we always learn new things

OpenStudy (lgbasallote):

@hartnn no

OpenStudy (lgbasallote):

wait... it seems i have my notations wrong...

OpenStudy (anonymous):

I think so. Shouldn't it be:  =1/3?

OpenStudy (lgbasallote):

\[\frac 13 = \cdot 3 \cdot \cdot\]

OpenStudy (lgbasallote):

anyway...you can use the cdots anyway you want... either for decimal point or ellipsis

OpenStudy (lgbasallote):

but only for decimals and ellipsis... ellipsis are used to denote repeating decimals...

OpenStudy (anonymous):

I would agree with: \[\frac{1}{3}=0.3...\] ... is an elipsis. \[\cdot \cdot \cdot\] (using cdots) is not.

OpenStudy (amistre64):

\[\overbrace {.345}\] lol, "bar" only works on one digit, the overbrace seems to span them ... just looks wierder

OpenStudy (lgbasallote):

the notation doesn't really matter on the solution does it?

OpenStudy (amistre64):

the rules are still a bit vague to me

OpenStudy (anonymous):

The notation is confusing. You said that: \[\cdot3=\frac{1}{3}\] So what does: \[\cdot0=?\] Just 0? I think you are confusing notations.

OpenStudy (unklerhaukus):

Lgba , please do not make up notation, i do not understand this question

OpenStudy (lgbasallote):

it's not a made up notation

OpenStudy (lgbasallote):

you may use the dots either as decimals or ellipsis

OpenStudy (unklerhaukus):

well i dont like it ,

OpenStudy (lgbasallote):

ellipsis would denote repeating decimals

OpenStudy (unklerhaukus):

\[\frac 19=0.111...=0.\dot1\] \[\frac1{99}=0.010101...=0.\dot0\dot1\]

OpenStudy (anonymous):

I don't think ellipsis is a good way of denoting repeating decimals. For example you could have: \[0.14159...\] Now, does that mean 0.141599999999999999, or 0.141591415914159?

OpenStudy (lgbasallote):

...why the heck are people meticulous about notations?? it doesn't really matter in the problem...

OpenStudy (lgbasallote):

this riddle is actually much easier than the original...since in the original, there's no explanation for what those dots mean

OpenStudy (lgbasallote):

i already explained what the dots mean...

OpenStudy (anonymous):

Notation is hugely important in mathematics. Just look at the argument between Newton and Leibniz over their simultaneous and independent discoveries of differential calculus. Leibniz' notation is now standard because it is so simple and easy to use, whereas Newton's only really finds use in mechanics and a few other places.

OpenStudy (lgbasallote):

it's a riddle...not a math quiz...in riddles you don;t stick to rules

OpenStudy (lgbasallote):

you go out-of-the-box

OpenStudy (unklerhaukus):

i still do not understand the question,

hartnn (hartnn):

if u post such confusing riddles, everyone instead of going out-of-the-box , will go out-of-the-post....

OpenStudy (anonymous):

If you expect people to attempt it then you should probably at least make the effort to make the problem well defined.

OpenStudy (lgbasallote):

that's the point

OpenStudy (lgbasallote):

and like i said @hartnn it is much harder because the original riddle didn't have clues like i gave

OpenStudy (lgbasallote):

it just gave .4.5.6.7.8.9.0.

OpenStudy (lgbasallote):

so it was A LOT more confusing

OpenStudy (lgbasallote):

so is that not well defined?

OpenStudy (anonymous):

Then you would have no way of knowing what the operation: \[\cdot\] defines. So I could just say that: \[\cdot(a, b, c, ..., z) = 82\] By defining the operation \cdot myself.

OpenStudy (unklerhaukus):

is the sum, a sum of seven terms?

OpenStudy (lgbasallote):

but i already gave a hint how to solve it

OpenStudy (lgbasallote):

like i said... just use the dots as decimals or ellipsis

hartnn (hartnn):

and since u are asking sum, we can use + in between right ?

OpenStudy (lgbasallote):

and it's a sum...but not guarranteed of 7 terms

hartnn (hartnn):

ok, so the solution is something like 80.5+.74+.6+ something like that ?? then we can atleast attempt.....

OpenStudy (lgbasallote):

yes something like that

hartnn (hartnn):

i think i m close ... its 80+(something)=82 now i just need to play with those dots and remaining numbers to get something = 2